# An LLM-based Readability Measurement for Unit Tests’ Context-aware Inputs

### v1

@@ 5.3 RQ3: Programmers’ Readability Measurements
In this RQ, we aim to determine whether programmers agree with C3’s measurement on tests’ readability.
5.3.1 Methodology
@@
(2) Questionnaire. We arrange these ten pairs into a questionnaire, forming ten questions by applying the question framework depicted in Fig. 7 to each pair of tests. We shuffle a pair’s tests to ensure that a manual or an automated test is not always shown in the Test 1 or Test 2 of the framework.
(3) Participants. Following the previous study [13], we invite $30$ participants to answer the questionnaire. Table IX shows their jobs and programming experiences.
TABLE IX: Participants’ jobs and programming experiences
@@ 5.3.2 Results
TABLE X: Participants’ choice summary
@@
(1) Choice summary. Table X summarizes participants’ choices. When C3 asserts the manual test is more readable, 88 out of 150 instances ($58.7\%$) show agreement, while 47 instances ($31.3\%$) display explicit disagreement. For the C3-Equal cases, explicit preferences are expressed in $52+88=140$ instances ($93.3\%$).
@@ 5.3.3 Discussion
(3) Unreadable inputs are necessary. Two participants mentioned that unreadable inputs are necessary when testing boundary conditions since many defects are triggered by unexpected inputs. This posed a challenge for C3 to measure tests since C3 cannot identify unreadable inputs that are on purpose. Still, our experimental setup avoided this issue: For a context, as long as one input meets the requirement, we assert that context is to be satisfied (see Sec. 5.2).
Answer to RQ3 When C3 identifies readable differences between tests, programmers tend to reach a consensus and agree with it, although C3 alone does not include all readability factors.
@@ 6 Threats to Validity
Construct Validity. In Sec. 3.1, we use NLP’s named entities [48, 49, 39, 50] to construct string type’ contexts and number formats to construct number types’ contexts, which brings a threat that such named entities are not complete to represent all contexts of real-world projects. We provide an example in Sec. 5.1. Firstly, We plan to extend C3 to mine contexts beyond prepared contexts in the source code automatically. Secondly, in practice, we allow C3’s users to integrate new contexts into C3 by providing their names, descriptions, and some examples for LLMs.
@@ 7 Conclusion
This paper proposes the Context Consistency Criterion (C3), an automatic readability measurement tool for primitive-type test input. C3 can mine the readability contexts for primitive-type parameters in the code under test and use these contexts to judge tests’ consistency as a readability metric. Besides, this paper proposes EvoSuiteC3 (i.e., EvoSuite enhanced by C3). Our experimental results show the distribution of readability contexts and confirm C3’s accuracy. Furthermore, while LLM-based test generation and EvoSuiteC3 surpass manual tests under C3 measurement, two traditional tools (Randoop and the original EvoSuite) almost fail to generate readable string-type inputs.
@@ 8 Data Availability
C3’s implementation and the experimental data are at https://doi.org/10.5281/zenodo.13028768.

### v2

@@ 5.4.1 Methodology
(2) Questionnaire. We arrange these ten pairs into a questionnaire, forming ten questions by applying the question framework depicted in Fig. 7 to each pair of tests. We shuffle a pair’s tests to ensure that a manual or an automated test is not always shown in the Test 1 or Test 2 of the framework. Note that we aim to determine whether C3 reflects programmers’ requirements for test readability. Hence, the questionnaire only asked interviewees which test was more readable because we did not want to inform respondents explicitly about differences in test inputs. Otherwise, it could cause respondents to focus too much on the inputs, influencing their judgments about the tests’ readability. After completing the questionnaire, we actively communicated with the respondents about the impact of the inputs on readability.
(3) Participants. Following the previous study [14], we invite $30$ participants to answer the questionnaire. Table X shows their jobs and programming experiences.
TABLE X: Participants’ jobs and programming experiences
@@ 5.4.2 Results
TABLE XI: Participants’ choice summary
@@
(1) Choice summary. Table XI summarizes participants’ choices. When C3 asserts the manual test is more readable, 88 out of 150 instances ($58.7\%$) show agreement, while 47 instances ($31.3\%$) display explicit disagreement. For the C3-Equal cases, explicit preferences are expressed in $52+88=140$ instances ($93.3\%$).
@@ 5.4.3 Discussion
(3) Unreadable inputs are necessary. Two participants mentioned that unreadable inputs are necessary when testing boundary conditions since many defects are triggered by unexpected inputs. This posed a challenge for C3 to measure tests since C3 could not identify unreadable inputs on purpose. Still, our experimental setup avoided this issue: For a context, as long as one input meets the requirement, we assert that context is to be satisfied (see Sec. 5.2).
Answer to RQ4 When C3 identifies readable differences between tests, programmers tend to reach a consensus and agree with it, although C3 alone does not include all readability factors.
6 Case Study: Adding a new Context
In C3’s Mining Component (Sec. 3.1), we prepare a group of contexts to restrict LLMs to choosing one of them or answering no context for parameters. Such design choice prevents LLMs from answering open-ended questions, thus mitigating LLMs’ hallucinations [48, 49]. However, our prepared contexts might not cover all the contexts of real-world projects. To cope with this issue, we allow users to add new contexts to C3. We leverage an example for our evaluation (Sec. 5.1) to show how to add new contexts.
This example comes from OpenMRS [64] (an application recording medical events). It contains a string parameter, role. This parameter requires readable inputs, and some instances include “Medical Student”, “Doctor”, “Data Manager”, etc. However, it was not recognized with a context in our evaluation since we did not define MRS’s Role context in Sec. 3.1. After adding the context name (“MRS_ROLE”) and some input examples like the above-mentioned instances, C3’s prompt for LLMs to mine contexts (see Fig. 3) added one more option for LLMs’ response. We used a concrete LLM (GPT-4-Turbo) to test C3, and C3 successfully recognized this parameter with the ”MRS_ROLE“ context.
Meanwhile, we prepared two tests shown in Code 4. Using GPT-4-Turbo, C3’s Judging Component (Sec. 3.2) also gave the correct result that “Nurse” is readable while “xxx” is not. “Nurse” is not one of the predefined input examples, confirming LLMs’ flexibility in the judging task. Note that in Sec. 3.2, we also use traditional NLP tools (e.g., CoreNLP [39]) to judge input readability via the named entity recognition (NER) function. However, such tools could not be applied to customized contexts like “MRS_ROLE” without manual extension. Conversely, LLMs’ support for arbitrary inputs allows us to support custom contexts easily.
1void test1() {
2 Role role = new Role("xxx");
3}
4void test2() {
5 Role role = new Role("Nurse");
6}
Code 4: Two tests for the role parameter
@@ 7 Threats to Validity
Construct Validity. In Sec. 3.1, we use $23$ named entities [51, 52, 39, 53] to construct string type’ contexts, which brings a threat that such named entities are not complete to represent all contexts of real-world projects. We provide an example in Sec. 5.1. Firstly, We plan to extend C3 to mine contexts beyond prepared contexts in the source code automatically. Secondly, in practice, we allow C3’s users to integrate new contexts into C3 (see Sec. 6).
@@ 8 Conclusion
This paper proposes the Context Consistency Criterion (C3), an automatic readability measurement tool for primitive-type test input. C3 mines the readability contexts for primitive-type parameters in the code under test and uses these contexts to judge tests’ consistency as a readability metric. Besides, this paper proposes EvoSuiteC3 (i.e., EvoSuite enhanced by C3). Our experimental results show the distribution of readability contexts and confirm C3’s accuracy. Furthermore, while LLM-based test generation and EvoSuiteC3 surpass manual tests under C3 measurement, two traditional tools (Randoop and the original EvoSuite) almost fail to generate readable string-type inputs.
@@ 9 Data Availability
C3’s implementation and the experimental data are at https://doi.org/10.5281/zenodo.13254261.
Acknowledgements
This paper is partially sponsored by the Shanghai Sailing Program (No. 22YF1428600) and the National Natural Science Foundation of China (No. 62202306).